# FonTS: Text Rendering with Typography and Style Controls

### v1

@@ 3.2 Style Control Adapters
Txt-Attn(Ours) 5.10% 81.43 55.00
Txt+Img-Attn 9.40% 77.92 31.00
Table 3: Ablation studies on the selection of parameters during TC-finetuning. Assessing the basic text rendering capabilities with different trained components of the DiT backbone.
@@
Ours w/ TC-FT 54.29 63.75
Ours w/o TC-FT 41.90 16.50
Table 4: Ablation studies examining the impact of TC-finetuning (TC-FT) on artistic text rendering.
Figure 7: Qualitative results on the style consistency in artistic text rendering compared with Flux [1] and SD3 [12].
4 Experiments
Text Rendering Benchmark. To assess the text rendering capabilities with word-level typography and style controls, we extend the existing scene text rendering benchmark [8] by introducing new benchmarks for basic text and artistic text rendering. BTR-bench. To evaluate word-level typography controls in basic text rendering, we introduce the Basic Text Rendering benchmark (BTR-bench). BTR-bench includes 100 prompts of different fonts and typographic attributes. For each text prompt, typographic attributes are randomly applied to three positions within the text to assess the model’s ability to render specific typographic attributes on individual words, while font attributes are applied to the entire text in the image. ATR-bench. To evaluate artistic text rendering, we introduce the Artistic Text Rendering benchmark (ATR-bench). Based on the single-letter and multi-letter classification in [36], we categorize the content into single-word and multi-word groups. Drawing on the style prompts from the GenerativeFont benchmark [25], we generate artistic individual letters and words using Flux [1]. These generated artistic letters and words are used for single-word and multi-word text rendering, respectively.
Style Control Dataset. Style control training consists of two phases, each utilizing different datasets. The first phase involves pretraining with a dataset of general image-text pairs. The second phase is fine-tuning, using a dataset that includes artistic text images and paired descriptions. For pretraining, we assembled a dataset called SC-general, which includes approximately 580k general image-text pairs with high aesthetic scores. These images were sourced from open-source datasets [11, 34]. For the fine-tuning phase, we created the SC-artext dataset. We compile a list of style descriptions and a list of words. Combining these lists generated various prompts for artistic text images, which were then used as input for Flux.1-dev [3], resulting in approximately 20k high-quality images. To ensure the images matched the original text content, we used shareGPT4v [10] to regenerate captions. Details of these two datasets can be found in the supplementary material.
Implementation Details. We use Flux.1-dev [1] as the base model. In TC-finetuning, we omit the text condition from the CLIP text encoder, as the pooled text embedding is not fed into the joint attention of the DiT backbone. For training the style control adapters, we select CLIP [29] as the image encoder since its visual embeddings are not sensitive to text [20], ensuring minimal impact on the content of the text while modifying the style.
Training details. For TC-finetuning, we fine-tune the base model for 40k steps using the TC-Dataset, incorporating a regularization prefix (‘sks’) in the text prompts. The batch size is set to 2. For the style control adapters, we train for 100k steps on the SC-general dataset and 15k steps on the SC-artext dataset, using a batch size of 4. All training is conducted on 8 A100 GPUs, with a learning rate of $1\times 10^{-5}$. To optimize memory usage, we backpropagate the loss every two steps and utilize DeepSpeed ZeRO [30].
4.1 Quantitative Results
For quantitative evaluation, we use the same OCR tool [4] as in [37] to detect text content in the generated images and calculate the OCR accuracy (OCR-Acc) across three distinct text rendering tasks. In the basic text rendering (BTR) task, existing OCR tools struggle to evaluate word-level typographic attribute accuracy (Word-Acc). Therefore, we use GPT4o and manual screening to assess and obtain the corresponding score. Font consistency (Font-Con) in BTR and style consistency (Style-Con) in artistic text rendering (ATR) are evaluated through user studies. Additionally, we compare our method to Flux.1 in the scene text rendering (STR) task (using the MARIO-bench [8]), evaluating OCR-Acc and CLIP scores (CLIP). Since Glyph-ByT5 and TextDiffuser-2 do not support the artistic text rendering defined in our study (we provide their qualitative results on the ATR-bench in the supplementary materials), the ATR column for these methods lacks quantitative results.
As shown in Table 1, our method outperformed the baselines in four out of five metrics while slightly below Glyph-ByT5 regarding OCR accuracy in BTR. This is reasonable since Glyph-ByT5 was trained on millions of text images, whereas our approach utilized a dataset of 50k basic text images, which is twenty times smaller.
Furthermore, as indicated in Table 2, our method significantly improves OCR-Acc and CLIP scores compared to Flux. Upon reviewing the image results, we found that Flux exhibits semantic confusion in the STR task (on MARIO-bench [8]), similar to its performance in ATR, which notably reduces its OCR accuracy.
User Studies. We conducted user studies with 22 participants to perceptually evaluate our results, comparing them against baseline methods. The evaluation included questions focused on two key aspects: font consistency (Font-Con) and style consistency (Style-Con). Font-Con was assessed through two subtypes: the first measured the consistency between the output image and the ground truth, while the second assessed font consistency across different outputs under identical font input conditions. Style-Con was evaluated similarly, with two subtypes covering the same aspects. The number of questions for each subtype is: 4, 2, 3, and 2, respectively. The score for each method was calculated based on the number of votes it received divided by the total number of votes. According to the results, our method achieved the best performance in the Font-Con and Style-Con categories, as shown in Table 1. Further details are provided in the supplementary materials.
4.2 Qualitative Results
Basic Text Rendering. We use a set of challenging prompt words for evaluation. For Flux.1 and Glyph-ByT5, the prompt example (for the leftmost images) is: “Blue Text: ‘Love knows no limits’ in Font: Josefin Sans, Add underline to ‘Love’, Background: pure yellow”. This prompt specifies the font applies a particular typographic attribute to a word and defines the text and background colors. As shown in Figure 5, Glyph-ByT5 achieves better font consistency than Flux but lacks word-level control. In contrast, our method ensures strong font consistency and enables word-level control, such as underline, bold, or italic.
Artistic Text Rendering. As TextDiffuser-2 and Glyph-ByT5 do not support artistic text rendering, We compare our method with SD3 and Flux. All qualitative results for our method are displayed with an image scale set to 0.9. We conduct solid experiments and have the following findings:
Semantic confusion. SD3 and Flux do not process image inputs, while ours is capable of handling both images and text prompts. To this end, we also report the results generated by Midjourney [23] using the same inputs as ours. As shown in Figure 6, in contrast to other approaches, our method effectively integrates style control without compromising the readability and integrity of the original text content. Meanwhile, it greatly mitigates the phenomenon of semantic confusion existing in all three comparison baselines.
Style consistency. To ensure that SD3 and Flux produce artistic text images with a consistent style, we set a fixed seed during generation and employed the same style prompt for all inputs, varying only the text content. Our method utilized the artistic single letter Flux generated with the same style prompt ( ‘A’ and ‘n’ in Figure 7), for style controls. The text prompt follows a simple format: Text: ‘content’ where content refers to the desired text for generation.
As depicted in Figure 7, results generated by SD3 have poor accuracy in some words, with errors in both letter transcription (e.g., ‘Banana’ becoming ‘BNANA’ in the second image of the first row) and case sensitivity (e.g., all uppercase letters in the sixth image of the first row). The middle row shows results from Flux, also revealing semantic confusion, notably for words with strong semantic like ‘parrot’, where the red parrot replaces the ‘p’ in ‘parrot’ as seen in the fifth image of the second row. The results in thelast row displays demonstrating superior text content accuracy and style consistency of our method compared to the baselines.
In addition, we find that the typography controls in BTR can be transferred to ATR and STR, as shown in  1 to a certain extent. It is reasonable that the degree of controllability will be affected by the given style image, particularly when it contains text. However, considering we only train the model using basic text images to learn those typography attributes, it shows considerable potential for generalization and domain adaption ability in the proposed method.
4.3 Ablation study
Ablation on TC-Finetuning. To assess the effectiveness of TC-finetuning, we performed ablation studies on the fine-tuned module, training two configurations: joint text attention (Txt-Attn) and joint text and image attention (Txt+Img-Attn). The results in Table 3 show the performance differences across these two settings for the BTR task. The data indicates that increasing the training parameters for the Img-Attn component actually reduces model performance on the BTR task, suggesting that fine-tuning only the Txt-Attn is a more effective approach. Meanwhile, we investigated the impact of TC-finetuning on artistic text rendering. By comparing the performance before and after finetuning, we observed a significant improvement in OCR accuracy, particularly for the multi-word subset, as shown in Table 5.
Ablation on ETC-Tokens. The ETC tokens are designed for assigning the words which need to be affected. To this end, assume we want to make the word ‘Find’ Bold of ‘Find your path’, we consider three cases: 1) no tokens: generate the results using language to express the intentions, i.e. using the prompt as ‘make the Find Bold in the content of Find your path’; 2) single token: following previous research [16, 6], we trained the model use single token and put the modifier token in front of the ‘Find’; 3) Ours. It can be found that the generated results using the first way brings great semantic confusion and the second way generated random results instead of accurately affecting the specific words. Details present in the supplementary materials.
@@ 4.4 Applications and Limitation
Stylization of scene text image. Since visual embeddings of CLIP [29] are not highly sensitive to text [20], using CLIP as the image encoder for style control adapters has minimal impact on the content of the text. Instead, it primarily affects the stylistic presentation. As shown in Figure 9, after injecting different style prompts, the style of text changes, but the content of text remains clearly identifiable.
@@
Limitation. It is observed the language drift phenomenon exists in our method, as the same as [16, 32]. This effect becomes noticeable as the number of training steps increases. This is mainly because, in the TC-finetuning process, we did not use additional regularization datasets; instead, we applied a simple regularization prefix, ‘sks’, in the text prompts of the TC dataset. This way decreases the cost. As shown in Figure 11, although language drift is severe at 60k steps, leading to the separation of text and scene in the generated image, the results at 40k steps are acceptable.
@@
5 Conclusion and Future Work
This paper proposes a two-stage DiT-based pipeline for text rendering with typography and style controls. TC-finetuning with ETC-tokens enables the model to learn and apply word-level typographic attributes to specific words. Style Control Adapter allows for independent style control using images. Experimental results show that the proposed method outperforms baselines on font consistency and word-level control in basic text rendering (BTR) and style consistency in artistic text rendering (ATR). To our knowledge, this paper is the first to achieve word-level control in basic text rendering, offering a valuable reference for future research in both scene and artistic text rendering. In future work, we aim to refine the scalability of our method further to handle a broader range of typography controls and styles and explore its extension to multilingual text rendering.
@@ 5 Conclusion and Future Work
{strip}[Uncaptioned image]
@@
Figure 14: Examples of images in SC-general.
Figure 15: Examples of SC-artext. All these artistic text images are generated by Flux.1-dev [1].
Figure 17: Examples of Typographic Controls in Scene Text Rendering.
@@
This supplementary material serves as a complement to the main paper, including demonstration of BTR, ATR and STR in Section A; details of the datasets used in Section B; details of user study in Section C; additional results presented in Section D; and further details regarding word accuracy (Word-Acc) in Section E.
Appendix A Demonstration of BTR, ATR and STR
@@
Appendix B Details of Datasets
This section complements Section 3 , 4 of the main paper, detailing the datasets we utilized in our study.
Typography Control Dataset (TC-Dataset). To address the lack of high-quality datasets that integrate text with word-level typographic attributes, we developed the TC-Dataset using typography control rendering (TC-Render). This process harnesses HTML rendering to generate images that display typographic features such as various fonts and word-level attributes, including bold, italic and underline. We initiated our process by extracting 625 text excerpts from novels. For each excerpt, we designed an HTML structure comprising sixteen images: one without typographic attributes and, in five different positions, applied three distinct typographic attributes (shown in Figure 13 (a)). Furthermore, we applied data augmentation techniques by randomly altering the text color and background (shown in Figure 13 (b)). Each HTML structure was rendered with one of five different fonts, resulting in approximately 50k text-image pairs set against solid color backgrounds.
@@ Appendix B Details of Datasets
SC-general. To train our style control adapters, we assembled the SC-general dataset, which includes approximately 580k general image-text pairs with high aesthetic scores. These pairs were sourced from open-source datasets [11, 34]. Figure 14 presents sample images, and Table 6 displays the corresponding paired texts.
SC-artext. For fine-tuning the style control adapters, we created the SC-artext dataset. We combined a list of 100 style descriptions with a list of 99 words, categorized into three character length groups: 1-15, 16-30, and 30-50. This combination produced a variety of prompts for artistic text images, which served as input for Flux.1-dev [1], yielding around 20k high-quality images. To ensure the images accurately reflected the original text content, we utilized shareGPT4v [10] to regenerate captions. Figure 15 shows sample images, and Table 6 presents the paired texts.
Appendix C Details of User Study
This section complements Section 4.1 of the main paper, providing additional details on the user studies. We involved 22 participants in these studies to evaluate our results perceptually, comparing them to baseline methods. The evaluation focused on two main aspects: font consistency (Font-Con) and style consistency (Style-Con). Style-Con was evaluated in two ways: one subtype measured the consistency between the output image and the ground truth, while the other assessed the consistency of fonts across different outputs when the same font input was used. This can be seen in Questions 1 and 2 in Figure 22. Font-Con was evaluated in a similar manner, with two subtypes addressing the same two aspects. These are represented by Question 3 of Figure 22 and Question 4 of Figure 23. Each subtype had a different number of questions: 4, 2, 3, and 2. The score for each method was determined by dividing the number of votes it received by the total number of votes cast.
Appendix D More Results
D.1 Ablation on ETC-Tokens
This section supplements Section 4.3 of the main paper, focusing on demonstrating the effectiveness of the proposed Enclosing Typography Control (ETC)-tokens for targeted word-level typographic attributes. For instance, to bold the word ”robot” in the phrase “i am not a robot”, we explore three settings: 1) Non-Token: Using a instruction prompt instead of adding modifier tokens, such as “Make ‘robot’ bold”. 2) Single-Token: Drawing on prior research [16, 6], we trained our model to use a single token, placing the modifier token before “robot”. 3) Our ETC-Token. The quantitative comparative results are presented in Table 5.
[…]
Non-Token 71.00 25.00
Single-Token 72.88 32.00
ETC-Token(Ours) 81.43 55.00
Table 5: Ablation studies of ETC-Token on basic text rendering.
D.2 Typographic Controls in STR
D.3 More Qualitative Results of ATR
This section supplements Section 4.1 of the main paper, offering a qualitative comparison of our method with Glyph-ByT5 [20] and Textdiffuser-2 [9] on the ATR-bench dataset, as depicted in Figure 19. Additionally, our extended qualitative results on the ATR-bench dataset, including single-word and multi-word examples, are presented in Figure 20.
D.4 Semantic Confusion
Appendix E Details about Word-Acc

### v2

@@ 4.3 Ablation study
Tokens only 0.78% 66.52
T5 text encoder 28.23% 0
Txt-Attn (Ours) 5.03% 82.85 55.00
Txt+Img-Attn 9.29% 77.92 31.00
Table 5: Ablation of different modules during TC-FT on BTR.
@@
60.07 28.89 44.48
58.09 65.39 61.74
65.12 34.48 49.80
64.27 60.07 62.17
Table 6: Ablation studies of fine-tuning with SC-artext (Art-FT) for SCA (on MM-DiT and Single-DiT both) and typography control fine-tuning (TC-FT) for backbone. The last row is ours.
Ablation on SCA. Our style control adapters (SCA) are trained through two phases as mentioned in Sec  3.2, and ablation is focus on the second phase. Comparing top two rows in Figure 7, it is evident that TC-FT enhances text accuracy, yet severely weakens the artistry. Shifting to the third row, Art-FT significantly boosts artistry without damaging the accuracy. The fourth row, being nearly identical to the third, suggests that after Art-FT, TC-FT have a minimal negative impact on artistry. Results presented in Table 7 further validate this observation within the ATR task, and we also compared SCA on MM-DiT only with SCA on MM-DiT and Single-DiT both. The results are detailed in supplementary Sec 2.1.
Ablation on ETC-Tokens. The ETC tokens are designed for assigning the words which need to be controlled. We consider three cases: 1) Non-token: directly use the prompt as “make the ‘word’ Bold”; 2) single token: use single token in front of the ‘word’ same in  [19, 7]; 3) Ours. The results are detailed in supplementary Sec 2.3.
@@ 4.4 Applications and Limitation
Limitation. It is observed the language drift phenomenon exists in our method, as the same as [19, 36]. This effect becomes noticeable as the number of training steps increases. This is mainly because, in the TC-FT process, we did not use additional regularization datasets; instead, we applied a simple regularization prefix, ‘sks’, in the text prompts of the TC dataset. This way decreases the cost. As shown in Figure 10, although language drift is severe at 60k steps, leading to the separation of text and scene in the generated image, the results at 40k steps are acceptable.
@@
5 Conclusion
We propose a two-stage DiT-based pipeline for text rendering with typography and style controls. TC-FT with ETC-tokens enables the model to learn and apply word-level attributes. The style control adapter facilitates style control without compromising text content. Additionally, we introduce the first word-level control dataset. Experimental results demonstrate that our method outperforms baselines in font consistency and style consistency, and word-level controls for text rendering tasks. This paper is the first to achieve word-level control in text rendering, in future work, we plan to explore its extension to multilingual rendering.
@@ 5 Conclusion
Figure 11: Examples of typographic controls in STR.
Figure 12: Examples of font selection in STR.
Figure 13: Ablation study of SCA only on MM-DiT with Art-FT and TC-FT.
Figure 14: Ablation study of SCA on MM-DiT and Single-DiT both, with Art-FT and TC-FT when image scale = 0.6.
Figure 15: Results of different backbones for the ID extractor in AnyDoor [12]. “DINOv2*” refers to removing the background of the target object with a frozen segmentation model before feeding it into the DINOv2 model. This figure is adapted from [12].
Figure 17: More qualitative results of ours on artistic text rendering.
Figure 18: Qualitative results of AnyText [41] and with TC-Finetuned on BTR.
Figure 19: The visualization of attention map on each word in different base models.
Figure 21: Visual results of ablation on ETC-tokens.
Figure 22: Ablation study of style control adapter (SCA), results from style captions only after 10k and 40k steps of TC-finetunig.
Figure 24: Semantic confusion can also be observed in SD3, Flux and Midjourney.
Figure 25: Results of fine-tuing T5 text encoder with new tokens, while input for CLIP is fix prompt: ‘words only, clean background’.
@@
Figure 28: Examples of images in SC-dataset, (a) is SC-general, and (b) is SC-artext.
@@
This supplementary material serves as a complement to the main paper, including additional results presented in Section A; more ablation studies of TC-FT, ETC-tokens, and SCA detailed in Section B; demonstration of BTR, ATR and STR in Section C; discussion of semantic confusion is detailed in Section D; details of the datasets used in Section E; details of word accuracy (Word-Acc) in Section F; and further details regarding user study in Section G.
Appendix A More Results
A.1 Typographic Controls in STR
A.2 Differences with Flux-IPA
1) Our style control adapters (SCA) employ a two-stage training approach. Fine-tuning with SC-artext significantly boosts artistry without compromising the accuracy of text, making it more suitable for ATR task.
2) In contrast to Flux-IPA(XLabs)  [^5] Flux-IPA(XLabs), which is only applied on MM-DiT, our SCA is implemented on both MM-DiT and Single-DiT to enhance style control, as depicted in Figure 13 with Figure  14. Even with a style image scale of 0.6, the style achieved by applying SCA on both MM-DiT and Single-DiT is markedly superior to that of applying SCA only on MM-DiT with a style image scale of 0.9. The comparison between Table 7 and Table 8 further validates this, as applying SCA on both on MM-DiT and Single-DiT yields a higher CLIP-I score under different settings.
3) Unlike Flux-IPA(InstantX)  [^6] Flux-IPA(InstantX) which uses SigLIP [50], our method select CLIP [33] as the image encoder. This choice is grounded in the distinct characteristics of these two models. SigLIP [50, 40] is renowned for its robust OCR capabilities. Conversely, as discussed in [23, 12], CLIP’s visual embeddings are insensitive to text. This insensitivity to text in CLIP’s visual embeddings is pivotal for our application, as it mitigates content leakage from style images (artistic text images). The visual outcomes presented in Figure 15 provide empirical evidence support for our selection.
4) Distinct from previous methods, we insert adapters in an interval-skip manner (on layer 0,2,4…) to reduce costs. In terms of parameter usage, the parameters of adapters in Flux-IPA(InstantX) is approximately 2.85 times of ours, as demonstrated in Table 9.
A.3 More Qualitative Results of ATR
This section serves as a supplement to Section 4.2 of the main paper, offering a qualitative comparison of our method with Glyph-ByT5 [23] and Textdiffuser-2 [10] on the ATR-bench dataset, as depicted in Figure 16. Additionally, we present our extended qualitative results on the ATR-bench dataset, including single-word and multi-word examples, are presented in Figure 17. Notably, in the second row of results in Figure 17, the accurate mirror reflection of letters in the every results further substantiates the effectiveness of our SCA. This example showcases that our SCA can inject style while meticulously maintaining text accuracy, providing additional empirical evidence for the capabilities of our proposed approach in the ATR task.
A.4 Train Baseline
In addition to the aforementioned comparisons, we fine-tune another baseline, AnyText [41] on the TC-dataset using a method similar to TC-finetuning. The quantitative results are presented in Table 11, while the qualitative results are shown in Figure 18. These results clearly reveals that AnyText fails to acquire word-level controllability. The performance of Glyph-ByT5 and Textdiffuser-2 exhibits similar limitations. This may be attributed to the inherent restricted capabilities of the base models for text rendering. Figure 19 shows the attention maps of different base models for different words in basic text rendering.
A.5 Stylization of STR
With our SCA, the influence of style input on the text within the image is minimal, as clearly observable in Figure 20. When distinct style images are incorporated, a pronounced transformation in the text style ensues. Notwithstanding these changes in style, the integrity of the text content is maintained, remaining accurate and distinguishable.
Appendix B More Ablation
B.1 Ablation on SCA
SCA Only on MM-DiT. Upon comparing Figure 13 and Figure 14, it is observed that when SCA is implemented on both MM-DiT and Single-DiT, the degree of stylization achieved is substantially greater than when SCA is applied solely to MM-DiT. This holds true even when the scale of the style image is lower (images Figure 14) in the former case (left images in Figure 13). A comparison between Table 7 and Table 8 provides additional validation of this assertion when evaluated in the context of CLIP-I metrics.
SCA with Art-FT and TC-FT. The CLIP-I and OCR-Acc presented in Table 7 are the average figures obtained on ATR task when the scale of the style image is set at 0.9 and 0.6 respectively. Table 7 is identical to Table 6 in the main paper. These values are placed here to enable a more direct comparison with SCA only on MM-DiT (Table 8). It becomes evident that, irrespective of whether SCA, the impacts of Art-FT and TC-FT on the ATR task remain consistent: Art-FT enhances stylization, while TC-FT improves content accuracy. Additionally, as shown in Table 10, after Art-FT, the degree of style degradation caused by TC-FT is reduced. This highlights the distinct but complementary roles of Art-FT and TC-FT in optimizing both the stylistic and content-related aspects of the results.
Without SCA. As is evident from Figure 22, in the absence of SCA, even when a detailed style caption is employed to characterize the style, diverse text contents result in inconsistent styles under the same random seed. Moreover, through a comparison of the images in the two rows, it becomes apparent that TC-FT exerts a certain degrading impact on the artistic style imparted by the style caption.
B.2 Ablation on TC-FT
Regarding the ablation study of typography control fine-tuning (TC-FT), we configured four distinct training scenarios: (1) only new tokens, (2) T5 text encoder with new tokens, (3) joint text attention (Txt-Attn) with new tokens, and (4) joint text-image attention (Txt+Img-Attn) with new tokens. As previously established in [15, 23], text rendering performance is primarily governed by the text encoder architecture. To explore this, we attempted to fine-tune the T5 on the BTR dataset to enhance controllability in text rendering. However, this approach led to a substantial decline in text accuracy, with visual artifacts evident in the generated outputs. The visual results are documented in Figure 25.
B.3 Ablation on ETC-Tokens
This section supplements Section 4.3 of the main paper, focusing on demonstrating the effectiveness of the proposed Enclosing Typography Control (ETC)-tokens for targeted word-level typographic attributes. For instance, to bold the word ”robot” in the phrase “i am not a robot”, we explore three settings: 1) Non-Token: Using a instruction prompt instead of adding modifier tokens, such as “Make ‘robot’ bold”. 2) Single-Token: Following [19, 7], we trained our model to use a single token, placing the modifier token before “robot”. 3) Our ETC-Token. The visual results of ablation on ETC-tokens are presented in Figure 21.
[…]
60.07 28.89 44.48
58.09 65.39 61.74
65.12 34.48 49.80
64.27 60.07 62.17
Table 7: Ablation studies of fine-tuning with SC-artext (Art-FT) for SCA (on MM-DiT and Single-DiT both) and TC-finetuning (TC-FT) for backbone. The last row is ours.
[…]
54.19 24.32 39.26
51.64 60.79 56.22
58.14 17.89 38.02
56.40 58.27 57.34
Table 8: Ablation studies of fine-tuning with SC-artext (Art-FT) for SCA (only on MM-DiT) and TC-finetuning (TC-FT) for backbone.
Modules Non-Skip Skip (Ours)
Adapters 1434.45 M 503.38 M
Table 9: Parameter quantity comparison with Flux-IPA(InstantX).
$\Delta_{CLIP-I}$ w/o Art-FT w/ Art-FT
Both 1.98 ($60.07\to 58.09$) 0.85 ($65.12\to 64.27$)
Only 2.55 ($54.19\to 51.64$) 1.74 ($58.14\to 56.40$)
Table 10: Comparison of CLIP-I changes with and without Art-FT in two SCA settings after TC-FT. Both: SCA on MM-DiT and Single-DiT both, Only: SCA only on MM-DiT.
[…]
AnyText 43.78 3.67
AnyText +TC-FT 39.26 2.64
Ours 82.85 55.00 68.42
Table 11: Quantitive results of AnyText and with TC-FT on BTR.
[…]
Non-Token 71.00 25.00
Single-Token 72.88 32.00
ETC-Token(Ours) 82.85 55.00
Table 12: Ablation studies of ETC-Token on basic text rendering.
Appendix C Demonstration of BTR, ATR and STR
@@
Appendix D Semantic Confusion
Additionally, we conducted additional comparisons with Midjourney [26], Flux, and SD3 in Figure 24. Whereas original SD3 and Flux lack the capability to process image inputs, both our proposed approach and Midjourney demonstrate the ability to handle combined image-text prompts. The results presented in the figure highlight a critical observation: during artistic text rendering tasks, semantic ambiguity significantly impairs the model’s capacity to accurately render the specified word’s content. Instead, the model tends to generate visual representations corresponding to the word’s semantic reference rather than words itself its. This phenomenon underscores the challenges inherent in balancing stylization and content accuracy within artistic text rendering.
Appendix E Details of Datasets
This section complements Section 3, 4 of the main paper, detailing the datasets we utilized in our work.
Typography Control Dataset (TC-Dataset). To address the lack of high-quality datasets that integrate text with word-level typographic attributes, we developed the TC-Dataset using typography control rendering (TC-Render). This process harnesses HTML rendering to generate images that display typographic features such as various fonts and word-level attributes, including bold, italic and underline. We initiated our process by extracting 625 text excerpts from novels. For each excerpt, we designed an HTML structure comprising sixteen images: one without typographic attributes and, in five different positions, applied three distinct typographic attributes (shown in Figure 27 (a)). Furthermore, we applied data augmentation techniques by randomly altering the text color and background (shown in Figure 27 (b)). Each HTML structure was rendered with one of five different fonts, resulting in approximately 50k text-image pairs with solid color backgrounds.
@@ Appendix E Details of Datasets
SC-general. To train our style control adapters, we assembled the SC-general dataset, which includes approximately 580k general image-text pairs with high aesthetic scores. These pairs were sourced from open-source datasets [14, 37]. Figure 28 (a) presents sample images, and Table 13 displays the corresponding paired texts.
SC-artext. For fine-tuning the style control adapters, we created the SC-artext dataset. We combined a list of 100 style descriptions with a list of 99 words, categorized into three character length groups: 1-15, 16-30, and 30-50. This combination produced a variety of prompts for artistic text images, which served as input for Flux.1-dev [1], yielding around 20k high-quality images. To ensure the images accurately reflected the original text content, we utilized shareGPT4v [11] to regenerate captions. Figure 28 (b) shows sample images, and Table 13 presents the paired texts.
Appendix F Details about Word-Acc